# SA-MLP: Enhancing Point Cloud Classification with Efficient Addition and Shift Operations in MLP Architectures

### v1

@@ Abstract
This study addresses the computational inefficiencies in point cloud classification by introducing novel MLP-based architectures inspired by recent advances in CNN optimization. Traditional neural networks heavily rely on multiplication operations, which are computationally expensive. To tackle this, we propose Add-MLP and Shift-MLP, which replace multiplications with addition and shift operations, respectively, significantly enhancing computational efficiency. Building on this, we introduce SA-MLP, a hybrid model that intermixes alternately distributed shift and adder layers to replace MLP layers, maintaining the original number of layers without freezing shift layer weights. This design contrasts with the ShiftAddNet model from previous literature, which replaces convolutional layers with shift and adder layers, leading to a doubling of the number of layers and limited representational capacity due to frozen shift weights. Moreover, SA-MLP optimizes learning by setting distinct learning rates and optimizers specifically for the adder and shift layers, fully leveraging their complementary strengths. Extensive experiments demonstrate that while Add-MLP and Shift-MLP achieve competitive performance, SA-MLP significantly surpasses the multiplication-based baseline MLP model and achieves performance comparable to state-of-the-art MLP-based models. This study offers an efficient and effective solution for point cloud classification, balancing performance with computational efficiency.
keywords:
Point Cloud , Classification , Shift , Addition , Computational Efficiency
journal:
1 Introduction
The analysis of point cloud data has become an essential task in numerous applications, including autonomous driving [1, 2], robotics [3, 4], and virtual reality [5, 6]. Unlike traditional 2D images, point clouds present unique challenges due to their irregular and sparse nature, which complicates the direct application of conventional neural network (CNN) architectures. Recent advances in deep learning, such as multi-layer perceptrons (MLPs), CNNs, and Transformers, have shown promise in processing point cloud data effectively [7, 8, 9]. However, the computational inefficiencies inherent in these models, primarily due to their reliance on multiplication operations, present significant barriers to real-time applications and deployment on resource-constrained devices .
Traditional neural networks depend heavily on multiplication operations within their layers, leading to high computational costs. Multiplicative operations require more clock cycles and consume more energy compared to simpler operations like addition or bit shifting. This inefficiency becomes particularly pronounced in large-scale models or those intended for deployment on edge devices. As a result, the search for alternatives to multiplication in neural network layers has gained momentum, with recent studies exploring the replacement of multiplication with more efficient operations such as addition and shift [10, 11, 12]. Given the increasing demand for efficient computation, optimizing the operations within neural networks has become critical. In this context, the present study aims to develop novel MLP-based architectures that leverage addition and shift operations to replace multiplications, thereby enhancing computational efficiency while maintaining or even improving classification performance on point cloud data.
In recent years, several studies have focused on optimizing CNNs by reducing the reliance on multiplication operations. AdderNet [10], for instance, replaces multiplications with additions in CNNs, leading to a significant reduction in computational cost while maintaining competitive performance. Similarly, DeepShift [11] introduces shift operations to replace multiplications, further improving computational efficiency by leveraging the simplicity of bitwise shifts. Both of these approaches have demonstrated that complex operations can be replaced by simpler alternatives without compromising accuracy, paving the way for more efficient deep learning models. Moreover, models like ShiftAddNet [12], which combine shift and addition operations, improve efficiency at the expense of increased parameter counts and limited representational capacity. Specifically, the ShiftAddNet [12] architecture replaces convolutional layers with a combination of adder and shift layers, leading to a doubling of both the parameter count and the number of layers. Additionally, ShiftAddNet [12] freezes the weights of the shift layers, which can restrict the model’s learning potential and overall performance. These limitations highlight the need for a more balanced approach that can harness the benefits of both addition and shift operations without introducing such drawbacks.
While most of the work on operation-efficient neural networks has focused on CNNs for image processing, similar ideas can be applied to point cloud analysis models. Traditional approaches such as MLP-based architectures rely on shared MLP layers, which, despite their effectiveness, suffer from high computational costs due to their use of multiplication-heavy layers. The challenge lies in adapting the principles of addition and shift operations to MLPs, where the absence of convolutional operations requires different strategies for efficiency gains.
Building upon the insights from previous works, this study proposes a series of novel MLP-based architectures tailored for efficient point cloud classification. First, we introduce the Add-MLP model, which replaces the standard multiplication operations within MLP layers with addition operations based on the $L_{1}$ norm. By doing so, Add-MLP significantly reduces computational complexity while maintaining comparable performance to traditional MLPs. However, purely relying on additions, while computationally efficient, may limit the expressiveness of the model due to the inability to capture certain multiplicative transformations.
Next, the Shift-MLP model is introduced, which substitutes multiplication with bitwise shift operations. This approach leverages the minimal computational cost of shift operations, further enhancing the efficiency of the model. However, the expressiveness of shift operations is inherently limited to scaling by powers of two, which restricts the model’s ability to capture the full range of multiplicative mappings possible with floating-point multiplications. As a result, networks that rely solely on shift operations may struggle to approximate a broader range of transformations, limiting their overall performance in complex tasks.
The cornerstone of this study is the ShiftAdd-MLP (SA-MLP) architecture, which synergistically combines the strengths of both addition and shift operations. Unlike ShiftAddNet [12], which replaces convolutional layers and doubles the number of layers and parameters, SA-MLP focuses on replacing MLP layers with alternately distributed adder and shift layers. This design maintains the original number of layers and avoids increasing the parameter count. By combining these operations, SA-MLP effectively addresses the limitations of the Add-MLP and Shift-MLP models. The adder layers contribute flexibility and fine-grained feature manipulation, while the shift layers offer coarse-grained efficiency and computational reduction. Furthermore, in contrast to the ShiftAddNet [12] model, where shift layer weights are frozen, SA-MLP allows for the learning of all parameters, thus preserving the model’s representational capacity .
@@ 1 Introduction
A crucial innovation of SA-MLP is the application of different learning rates and optimizers for the adder and shift layers. This tailored optimization strategy ensures that each type of operation can contribute optimally to the learning process, enhancing the overall model performance without compromising efficiency. By decoupling the learning strategies for the adder and shift layers, SA-MLP fully exploits their complementary advantages, resulting in a model that offers both high accuracy and reduced computational overhead.
The main contributions of this study are as follows:
  • 1.
@@
  • 2.
    We propose SA-MLP, which integrates adder and shift layers within MLPs, maintaining the original number of layers and parameters while ensuring full participation of all parameters in the learning process.
  • 3.
@@
    We develop a tailored optimization strategy that applies distinct learning rates and optimizers for adder and shift layers, effectively leveraging their complementary strengths to enhance model performance.
  • 4.
@@
    Extensive experimental evaluations demonstrate that the proposed SA-MLP outperforms the traditional multiplication-based MLP model and achieves comparable performance to state-of-the-art MLP-based models, offering a more efficient and scalable solution for point cloud classification.
2 Related works
2.1 Point Cloud Analysis
The burgeoning field of point cloud analysis has witnessed several innovative approaches leveraging deep learning to extract meaningful features from unstructured point sets. A pioneering work in this domain is the introduction of PointNet [7], which demonstrated the feasibility of directly processing point clouds with a deep learning framework. This was further expanded upon by PointNet++ [13], enhancing feature learning through a hierarchical structure that respects the metric space of point clouds. Convolutional Neural Networks (CNNs) have also been adapted to the point cloud domain, with PointCNN [14] standing out as a method that generalizes the convolution operation to point sets through a novel $\mathcal{X}$-Transformation. A parallel development in this space is the exploration of continuous convolutional neural networks [15], which aim to learn parametric filters for point cloud data. Graph Neural Networks (GNNs) have emerged as a powerful tool for point cloud analysis, with Dynamic Graph CNN [16] introducing a dynamic graph construction approach that adapts to the local geometry of point clouds. Building upon this, Adaptive Graph Convolution [8] has been proposed to further refine the feature extraction process through an adaptive scheme. In a significant departure from traditional convolutional and graph-based methods, Transformer architectures have made their mark in point cloud analysis with the advent of Point Transformer [17]. This approach utilizes the self-attention mechanism to capture global dependencies within point sets. Subsequent advancements have led to Point Transformer V2 [18], which introduces grouped vector attention and partition-based pooling to enhance the model’s capability.
2.2 Multiplication-Free Neural Networks
The pursuit of computational efficiency in deep learning has led to the emergence of Multiplication-Free neural networks, a paradigm shift aimed at reducing the computational complexity of convolutional operations. In this context, several approaches have been explored to construct neural networks without the need for multiplication, a conventionally resource-intensive operation.
Works such as BNN [19], XNOR-Net [20] and approach [21] have ventured into binarizing weights or activations, significantly curtailing the computational load by enabling operations to be performed using simple sign determinations instead of floating-point multiplications. This approach not only streamlines the computational process but also compresses the model size, facilitating deployment on devices with limited resources. Shift operations, another avenue explored by SSL [22], Shift [23] and DeepShift [11], offer an alternative to multiplication by leveraging the hardware efficiency of bitwise operations. These operations, being inherently faster and less resource-demanding, provide a compelling solution for accelerating neural network implementations. The innovative approach of AdderNet [10] and subsequent works like AdderSR [24] and approach [25], which rely on additive operations, challenge the traditional reliance on multiplication in neural networks. By employing addition as the fundamental operation, these networks demonstrate that it is possible to achieve high performance with reduced computational overhead. Furthermore, the hybrid approach combining both shift and add operations, as seen in ShiftAddNet [12] and ShiftAddViT [26], has shown to be particularly effective. These models harness the strengths of both operations, offering a flexible trade-off between accuracy and efficiency.
In summary, while Multiplication-Free neural networks have made strides in efficiency, their application to point cloud analysis is relatively unexplored. This study breaks new ground by integrating shift and add operations into point cloud classification, demonstrating that these networks can achieve high accuracy with reduced computational costs.
3 Methodology
@@ 3.1 Baseline Model: Mul-MLP
3.2 Shift-Based Model: Shift-MLP
The Shift-MLP model proposed in this study introduces an innovative approach by replacing the traditional multiplication operations in MLP layers with bitwise shift operations, significantly improving computational efficiency without compromising model performance. The underlying principle of the shift layer involves quantizing the weights to powers of two, thereby transforming multiplication into efficient shift operations. In the forward propagation, each weight $W$ is first quantized using the formulas $S=\text{sign}(W)$ and $p=\text{round}(\log_{2}(|W|))$, resulting in a quantized weight:
@@ 3.2 Shift-Based Model: Shift-MLP
3.3 Addition-Based Model: Add-MLP
@@
where $Y$ is the output, $X$ is the input, and $W$ represents the corresponding weights. The use of the $L_{1}$ norm allows the network to measure relevance based on additive differences, which fundamentally reduces the computational complexity compared to traditional multiplication-based operations.
@@
3.4 Combined Model: ShiftAdd-MLP (SA-MLP)
The SA-MLP model introduced in this study builds on the concepts of Shift-MLP and Add-MLP, offering an advanced approach that fully leverages the complementary strengths of shift and add operations (see Fig. 4). Unlike ShiftAddNet [12], which operates in the image domain and doubles the number of layers by replacing each convolutional layer with a pair of Shift and Adder layers, SA-MLP integrates these operations in an interleaved manner within the MLP framework. The key innovation lies in replacing the traditional MLP layers in the baseline Mul-MLP model with a sequence of interleaved Shift and Adder layers. This design ensures that the number of layers and parameters remains consistent with the original model, avoiding the significant increase in complexity observed in ShiftAddNet [12].

### v2

# SA-MLP: A Low-Power Multiplication-Free Deep Network for 3D Point Cloud Classification in Resource-Constrained Environments

@@ Abstract
Point cloud classification plays a crucial role in the processing and analysis of data from 3D sensors such as LiDAR, which are commonly used in applications like autonomous vehicles, robotics, and environmental monitoring. However, traditional neural networks, which rely heavily on multiplication operations, often face challenges in terms of high computational costs and energy consumption. This study presents a novel family of efficient MLP-based architectures designed to improve the computational efficiency of point cloud classification tasks in sensor systems. The baseline model, Mul-MLP, utilizes conventional multiplication operations, while Add-MLP and Shift-MLP replace multiplications with addition and shift operations, respectively. These replacements leverage more sensor-friendly operations that can significantly reduce computational overhead, making them particularly suitable for resource-constrained sensor platforms. To further enhance performance, we propose SA-MLP, a hybrid architecture that alternates between shift and adder layers, preserving the network depth while optimizing computational efficiency. Unlike previous approaches such as ShiftAddNet, which increase the layer count and limit representational capacity by freezing shift weights, SA-MLP fully exploits the complementary advantages of shift and adder layers by employing distinct learning rates and optimizers. Experimental results show that Add-MLP and Shift-MLP achieve competitive performance compared to Mul-MLP, while SA-MLP surpasses the baseline, delivering results comparable to state-of-the-art MLP models in terms of both classification accuracy and computational efficiency. This work offers a promising, energy-efficient solution for sensor-driven applications requiring real-time point cloud classification, particularly in environments with limited computational resources.
Index Terms:
3D sensors, point cloud, classification, deep learning, computational efficiency, multiplication-free neural networks.
I Introduction
The analysis of point cloud data has become a cornerstone in various sensor-driven technological domains, such as autonomous driving [1, 2], robotics [3, 4], and virtual reality [5, 6]. Point clouds, which represent 3D spatial information through a collection of discrete points, are typically captured by LiDAR and other 3D sensors, providing rich data that is crucial for many applications. However, the inherently irregular and sparse nature of point clouds makes them challenging to process using traditional deep learning architectures originally designed for structured data like 2D images. As sensor systems become more prevalent in edge applications, real-time data processing is becoming increasingly important. This study focuses on improving the computational efficiency of point cloud classification algorithms for resource-constrained sensor platforms, which are often limited in terms of processing power and energy.
Over the years, advanced neural network architectures, such as multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers, have demonstrated significant potential in processing point cloud data [7, 8, 9, 10]. Despite their effectiveness, these architectures often rely heavily on multiplication operations, which are computationally intensive and energy-demanding. Such reliance presents critical obstacles for their adoption in real-time sensor applications, particularly in edge devices or embedded systems where computational resources are constrained. This inefficiency is a major barrier in applications like autonomous driving and robotics, where sensor data processing must occur in real-time with limited energy consumption.
The computational inefficiency of multiplication operations in neural networks stems from their high energy consumption and increased clock-cycle demands compared to simpler alternatives such as addition or bitwise shifting. These inefficiencies become particularly problematic when deploying models on edge devices that are integral parts of sensor systems. This has motivated recent research into operation-efficient neural networks that replace multiplication with computationally lighter operations. For instance, AdderNet [11] replaces multiplications with addition operations in CNNs, achieving significant reductions in computational cost while maintaining competitive performance. Similarly, DeepShift [12] introduces bitwise shift operations as substitutes for multiplication, leveraging the simplicity of shift operations to achieve enhanced efficiency. These approaches have demonstrated that it is possible to maintain the performance of neural networks while significantly reducing their computational demands, thus making them more suitable for sensor-based applications. However, architectures like ShiftAddNet [13], which combine shift and addition operations, introduce their own challenges. Specifically, ShiftAddNet replaces convolutional layers with adder and shift layers, doubling both the parameter count and the number of layers. Additionally, it freezes the weights of shift layers, limiting the representational capacity and learning potential of the model. These trade-offs underscore the need for sensor-optimized approaches that balance computational efficiency with learning flexibility and representational power.
While much of the existing work on operation-efficient neural networks has focused on CNNs for image processing, the principles of replacing multiplication with addition and shift operations can be extended to MLP-based architectures. MLPs are widely used in point cloud processing due to their ability to capture point-wise features and their simplicity compared to CNNs. However, the reliance on multiplication-heavy layers makes MLPs computationally expensive and less suitable for real-time sensor processing. Unlike CNNs, which primarily rely on convolutional operations to process structured data, MLPs are designed to handle unstructured data such as point clouds through point-wise feature extraction. However, the absence of convolutional mechanisms in MLPs necessitates careful adaptation of multiplication-free designs to ensure compatibility with the unique requirements of point cloud data. Consequently, there is a need for novel strategies to achieve computational efficiency in MLPs while maintaining high classification performance, particularly in the context of sensor data processing for resource-constrained systems.
In this study, we propose a series of MLP-based architectures that address the computational inefficiencies of traditional point cloud classification models, with a focus on making them suitable for sensor applications. The baseline model, Mul-MLP, adopts the conventional MLP design and serves as a benchmark for comparison. To reduce computational costs, we introduce Add-MLP and Shift-MLP, which replace the multiplication operations in Mul-MLP with addition and bitwise shift operations, respectively. Add-MLP leverages addition operations based on the $L_{1}$ norm, achieving a significant reduction in computational cost while maintaining performance comparable to traditional MLPs. However, the expressiveness of Add-MLP is limited, as addition operations may struggle to capture certain multiplicative transformations. Similarly, Shift-MLP enhances computational efficiency by replacing multiplications with bitwise shifts but is inherently constrained by the scaling limitations of shift operations, which are restricted to powers of two. As a result, models relying solely on either addition or shift operations may lack the flexibility needed for complex tasks.
To address these limitations, we propose SA-MLP, a hybrid architecture that combines the strengths of addition and shift operations. SA-MLP replaces multiplication-heavy MLP layers with alternately distributed adder and shift layers, preserving the original depth and parameter count of the network. Unlike previous models such as ShiftAddNet [13], SA-MLP allows all parameters, including those in shift layers, to learn during training, avoiding the constraints imposed by frozen weights. By integrating addition and shift operations, SA-MLP leverages the complementary strengths of these approaches: the adder layers provide fine-grained flexibility for feature manipulation, while the shift layers enhance computational efficiency. This balanced design ensures that the model achieves high representational capacity without incurring unnecessary computational overhead.
@@ I Introduction
A key innovation of SA-MLP is its tailored optimization strategy, which assigns distinct learning rates and optimizers to the adder and shift layers. This approach allows each type of operation to contribute optimally to the learning process, maximizing the complementary advantages of the two operations. By decoupling the optimization strategies, SA-MLP effectively combines computational efficiency with high accuracy, making it well-suited for sensor systems operating in real-time, where both energy consumption and performance are critical considerations.
Extensive experiments validate the effectiveness of the proposed architectures. Both Add-MLP and Shift-MLP achieve competitive performance compared to traditional MLPs while improving computational efficiency by leveraging addition and shift operations in place of multiplications. SA-MLP, in particular, surpasses the performance of the baseline Mul-MLP model and achieves results comparable to state-of-the-art MLP-based architectures. Additionally, the hybrid architecture offers a scalable and efficient solution for point cloud classification, providing a practical alternative for applications where computational efficiency and low energy consumption are paramount. This is especially true for sensor systems where resources such as power and processing capacity are limited, making it ideal for real-time deployment on edge devices like LiDAR-based systems in autonomous vehicles and robotic systems.
This study makes the following contributions:
@@
  • We propose SA-MLP, a hybrid architecture that combines adder and shift layers within MLPs, maintaining the original network depth and parameter count while allowing all parameters to learn during training.
@@
  • We develop a tailored optimization strategy with distinct learning rates and optimizers for adder and shift layers, effectively leveraging their complementary strengths.
@@
  • Experimental evaluations demonstrate that SA-MLP outperforms the traditional multiplication-based Mul-MLP model and achieves performance comparable to state-of-the-art MLP-based models, offering a more efficient and scalable solution for point cloud classification.
II Related Works
The field of 3D point cloud analysis has experienced significant advancements, particularly with the integration of deep learning techniques that effectively process unstructured data for tasks such as object recognition and scene understanding. Concurrently, the pursuit of computational efficiency has led to the development of multiplication-free neural networks, which aim to reduce computational complexity by substituting traditional multiplication operations with more efficient alternatives. The following subsections delve into these two pivotal areas, providing an overview of key methodologies and their contributions to the advancement of efficient 3D point cloud analysis.
II-A Point Cloud Analysis
The analysis of 3D point clouds has become a focal point in computer vision, especially for applications in sensor-based systems such as object recognition, scene understanding, and autonomous navigation. Point clouds, which are typically generated by sensors like LiDAR (Light Detection and Ranging), depth cameras, and stereo vision systems, provide rich spatial data that can capture the geometry of physical environments. These sensors are foundational in fields such as autonomous vehicles, robotics, and geospatial mapping. Early breakthroughs in point cloud processing leveraged deep learning techniques to handle the unstructured nature of point cloud data, marking a departure from traditional grid-based representations used in 2D image processing.
One of the seminal works in this area is PointNet [7], which demonstrated that deep neural networks could directly process point clouds without requiring structured grid-like representations, a significant departure from traditional methods. PointNet++ [14] extended this idea by introducing a hierarchical architecture that captures local and global geometric information, thus improving the model’s ability to learn from complex spatial structures.
Alongside these pioneering approaches, Convolutional Neural Networks (CNNs) have been adapted to process point cloud data. One notable example is PointCNN [8], which generalizes the convolution operation to point clouds through the introduction of the $\mathcal{X}$-Transformation, effectively transforming point clouds into a more structured form suitable for convolution. Another promising direction involves the exploration of continuous convolutional networks [15], which aim to learn parametric filters for point cloud data, enabling more flexible and efficient feature extraction.
Graph Neural Networks (GNNs) have also shown great potential in point cloud analysis. Dynamic Graph CNN (DGCNN) [16] introduced a dynamic graph construction method, which adapts to the local geometry of the point cloud, allowing the network to focus on relevant local structures. Building on this, Adaptive Graph Convolution [9] proposed an adaptive graph convolution scheme to further refine the feature extraction process by considering the varying local point cloud characteristics.
In contrast to these graph-based methods, Transformer-based architectures have gained traction in point cloud analysis, with Point Transformer [10] being a notable example. Utilizing the self-attention mechanism, Point Transformer models capture global dependencies between points, enabling more robust feature learning. Point Transformer V2 [17] extended this idea by incorporating grouped vector attention and partition-based pooling, which further enhances the model’s ability to model complex dependencies and structures in point clouds.
In summary, the integration of deep learning techniques with sensor-based point cloud data has led to significant advancements in the analysis of 3D data. However, while these methods have shown promising results in research settings, challenges remain in terms of real-time implementation and computational efficiency, especially in the context of embedded sensor systems and resource-constrained environments. The evolution of models like PointNet [7], PointCNN [8], and Point Transformer [10] reflects the growing sophistication of algorithms tailored to the unique characteristics of sensor data, with the potential to significantly impact fields such as autonomous navigation, robotics, and environmental sensing.
II-B Multiplication-Free Neural Networks
In parallel with the advancements in point cloud analysis, the increasing need for computational efficiency in sensor-based systems has driven the exploration of multiplication-free neural networks. These networks aim to reduce the computational burden typically associated with traditional multiplication operations, which are computationally expensive, by replacing them with simpler, faster, and energy-efficient operations. This is particularly crucial for embedded sensor systems, such as autonomous vehicles, drones, and IoT devices, where real-time data processing, low latency, and low power consumption are essential.
Early efforts in this direction focused on binarizing weights and activations, as seen in Binary Neural Networks (BNN)[18] and XNOR-Net[19]. These approaches replace floating-point multiplications with binary operations, significantly reducing both computational load and memory footprint. These approaches replace floating-point multiplications with binary operations, significantly reducing both the computational load and memory footprint, thus making them suitable for deployment in resource-constrained environments.
Shift-based operations also offer a promising alternative to multiplication. For instance, methods like SSL [20], Shift [21], and DeepShift [12] leverage bitwise shifts, which are highly efficient on hardware, to replace multiplication. These approaches capitalize on the efficiency of simple shift operations to accelerate neural network implementations. Another key advancement is the use of additive operations in networks, such as AdderNet [11], which replaces multiplications with additions. The AdderNet framework, along with subsequent works like AdderSR [22] and AdderFPGA [23], demonstrates that high performance can still be achieved with additive operations, challenging the traditional reliance on multiplication.
A more recent trend combines both shift and add operations in hybrid models. ShiftAddNet [13] and ShiftAddViT [24] are examples of architectures that integrate both operations, providing a flexible trade-off between computational efficiency and model accuracy. These hybrid models exploit the advantages of both shift and add operations, making them particularly effective in reducing the computational overhead while maintaining competitive performance.
Despite the significant advancements in multiplication-free architectures, their application to point cloud analysis remains relatively underexplored. Most existing work on multiplication-free networks has focused on tasks like image classification, which do not have the inherent complexities of 3D point cloud data. Point cloud analysis—especially in sensor-based systems that rely on LiDAR or depth sensors—presents unique challenges due to the high dimensionality and unstructured nature of the data. This study contributes to filling this gap by introducing shift and add operations for 3D point cloud classification. Through extensive experimentation, we demonstrate that these multiplication-free architectures not only reduce computational costs but also achieve competitive classification accuracy when compared to traditional models, thereby providing a promising solution for efficient point cloud analysis in resource-constrained environments.
III Methodology
@@ III-A Baseline Model: Mul-MLP
III-B Shift-Based Model: Shift-MLP
The Shift-MLP model proposed in this study, inspired by DeepShift [12], introduces an innovative approach by replacing the traditional multiplication operations in MLP layers with bitwise shift operations, significantly improving computational efficiency without compromising model performance. The underlying principle of the shift layer involves quantizing the weights to powers of two, thereby transforming multiplication into efficient shift operations. In the forward propagation, each weight $W$ is first quantized using the formulas $S=\text{sign}(W)$ and $p=\text{round}(\log_{2}(|W|))$, resulting in a quantized weight:
@@ III-B Shift-Based Model: Shift-MLP
III-C Addition-Based Model: Add-MLP
@@
where $Y$ is the output, $X$ is the input, and $W$ represents the corresponding weights. The use of the $L_{1}$ norm allows the network to measure relevance based on additive differences, which fundamentally reduces the computational costs compared to traditional multiplication-based operations.
@@
III-D Combined Model: ShiftAdd-MLP (SA-MLP)
The SA-MLP model introduced in this study builds on the concepts of Shift-MLP and Add-MLP, offering an advanced approach that fully leverages the complementary strengths of shift and add operations (see Fig. 4). Unlike ShiftAddNet [13], which operates in the image domain and doubles the number of layers by replacing each convolutional layer with a pair of Shift and Adder layers, SA-MLP integrates these operations in an interleaved manner within the MLP framework. The key innovation lies in replacing the traditional MLP layers in the baseline Mul-MLP model with a sequence of interleaved Shift and Adder layers. This design ensures that the number of layers and parameters remains consistent with the original model, avoiding the significant increase in computational costs observed in ShiftAddNet [13].